A car starts from rest travels on a straight road and its wheel rotates 500 rotations in 15 s. If the diameter of a car’s wheel is 60 cm, calculate the distance traveled by the car.


1
Expert's answer
2021-11-01T19:24:42-0400

Solution. After 500 revolutions, the distance traveled is


"d=\pi\times d\times 500"

where d=0.6m is diameter of a car’s wheel. As result get

"d=\pi\times 0.6\times 500 \approx" 942m

Answer. 942m.